NIH | National Cancer Institute | NCI Wiki  




Discussion items


Java 17
  • Migration to Java 17
    • Corretto Java 17 installed on Dev
    • Tomcat 9 partially installed on Dev - ticket in for configuration

Oracle OS
  • CTS2 having issues since OS upgrade
  • Meeting with Zihua Feb 27.
    • Stage not consistent with lower tiers
    • CTS2 on Stage given 8g memory. Only 1-2g on lower tiers
    • Would like to see a configuration map of all servers
    • Different teams handle upper tiers vs lower tiers

Email/Contact Us
  • Programmatic options
  • CRM to filter messages
    • Would be a service that sets up a queue of real issues
    • "High noise, low signal"
    • Cost? $5-30/user.  What is a user?
  • Inventory public facing pages for email address
    • Can we get rid of Contact Us page?
    • Send email to user to confirm before sending onwards to final target?
  • Source appears to be the Contact Us page, as it is being sent to both NCIThesaurus and NCICBIIT
    • A lot is coming from an .ru address
    • We have a captcha on that page
  • Jason discussed with Lori about possibility of dropping form.
    • Some resistance from Lyubov to dropping functionality
  • Check if we can drop the Contact Us form?  The page would exist but the form portion would be gone. 

Production NCI Terms
  • Still multiple issues on Prod
    • Scott creating Jira issues as items identified
  • L3cache issues?  Possibly filling up and not allowing processing by cores
    • Dev has larger cache and more CPUs
  • Possible security solution - throttle IPs that are overworking CPU
    • L5 load balancer
  • Tracy to send time frames for issues- by the hour

System migration
  • Prod in testing
  • Stage 2 - can we ask for top access to this?
  • DataQA after Prod
    • database on machine - migrate first?
    • Test performance of old loaders vs new database.
    • Other apps follow DB - determine if needs to be co-located
    • Does script need to be updated?
  • List of EVS URLs
  • Check with systems - ask for testing meeting or access. Jason to put in ticket


ehcache issues on Meta, similar to what we have seen previously on NCI Terms.  Turned out to need removal of lb jar from server.  Now working.

SumoLogic to DataDog
  • Queries being migrated by Systems
    • Connectors updated except for Drupal server - Zihua He working on config
  • Dialogs being built from scratch
  • S3 bucket for archive storage
    • Need live testing of DataDog interaction


Can add and remove users through Service Now.

Processing server - update lvg

Current lvg2010 uses Java 7.  This needs to be updated per security

  • Tracy working to update ProtegeKbQA, processing script.
  • Will need testing on processing server once ready

Action items